The Blue Mountain High School Drama Club is selling tickets to their spring musical. Adult tickets are $4 and student tickets are $1. A total of 285 tickets are sold for $765. How many of each type of ticket are sold?
A) 145 adult, 140 student
B) 120 adult, 165 student
C) 180 adult, 105 student
D) 160 adult, 125 student

Answer by stanbon(75887)

The Blue Mountain High School Drama Club is selling tickets to their spring musical. Adult tickets are $4 and student tickets are $1. A total of 285 tickets are sold for $765. How many of each type of ticket are sold?

-------
a + s = 285 tickets
4a + 1s = 765
----
Multiply thru the 1st equation by 4:
4a + 4s = 4*285
4a + s = 765
------
Subtract and solve for "s":
3s = 375
s = 125 (# of student tickets sold)
---
Solve for a:
a + s = 285
a = 160 (# of adult tickets sold)
=====================
Cheers,
Stan H.
==============

A) 145 adult, 140 student
B) 120 adult, 165 student
C) 180 adult, 105 student
D) 160 adult, 125 student